# Finding Quasars behind the Galactic Plane. IV. Candidate Selection from Chandra with Random Forest

### v1

@@ III Method
III.1 Random Forest Classification
We employ a RF classifier to distinguish quasars from stars in the CSC–WISE–GAIA sample. RF is an ensemble learning algorithm based on decision trees and bootstrap aggregation, originally introduced by Ho1995 and later formalized by Breiman2001. By combining multiple weak learners trained on bootstrapped subsets of the data and randomly selecting input features at each split, the RF algorithm effectively mitigates overfitting and performs robustly in high-dimensional parameter space.
The RF classifier is implemented using the scikit-learn [^2] https://scikit-learn.org/stable/index.html machine-learning library (scikit-learn). Each decision tree is trained independently, and the final classification is determined by a majority vote across all trees. For each source, the fraction of trees voting for the quasar class is interpreted as the quasar membership probability, providing a quantitative measure of classification confidence.
To separate quasars from stars, we adopt a total of 31 input features that incorporate multi-wavelength photometric, astrometric, and X-ray information. These features include X-ray fluxes and hardness ratios from CSC 2.1, optical and mid-infrared magnitudes and colors from Gaia DR3 and WISE, extinction-related quantities, and astrometric parameters such as proper motion and parallax significance. A complete list and description of the adopted features are provided in Table 3.
@@ III.1 Random Forest Classification
$\rm{HR}_{hm}$ ACIS hard (2.0-7.0 keV) - medium (1.2-2.0 keV) energy band hardness ratio
$\rm{HR}_{hs}$ ACIS hard (2.0-7.0 keV) - soft (0.5-1.2 keV) energy band hardness ratio
$\rm{HR}_{ms}$ ACIS medium (1.2-2.0 keV) - soft (0.5-1.2 keV) energy band hardness ratio
nh_gal Galactic $N_{H}$ column density in direction of source
$\log(f_{x})$ logarithm of X-ray flux in 0.5-7keV ($b$ band)
@@
$\log(f_{\rm BP})$ logarithm of $BP$ band flux
$\log(f_{\rm RP})$ logarithm of $RP$ band flux
@@
$RP-\mathrm{W1}$ $RP-\mathrm{W1}$ color
$G-\mathrm{W1}$ $G-\mathrm{W1}$ color
$G-\mathrm{W2}$ $G-\mathrm{W2}$ color
$\mathrm{W1}-\mathrm{W2}$ $\mathrm{W1}-\mathrm{W2}$ color
Missing values may arise from limited survey depth, nondetections, or incomplete coverage in specific bands. To allow uniform processing of incomplete data, all missing entries are assigned a sentinel value of 9999-9999, enabling the RF algorithm to handle such cases consistently during training and prediction.
The RF model includes several tunable hyperparameters, whose optimization is critical for achieving optimal classification performance. We focus on three key parameters: the number of trees in the ensemble ($n_{\mathrm{estimators}}$), the number of features considered at each split ($max_{\mathrm{features}}$), and the maximum depth of individual trees ($max_{\mathrm{depth}}$). These hyperparameters are optimized using a grid search with five-fold cross-validation. The optimal configuration yields $max_{\mathrm{features}}=\log_{2}31$ and $n_{\mathrm{estimators}}=41$, while no explicit upper limit on $max_{\mathrm{depth}}$ is required based on the validation results.
After hyperparameter optimization, the labeled dataset is randomly divided into a training subset and an independent validation subset with a ratio of 4:1. The RF classifier is trained on the training subset and evaluated on the validation subset. Classification performance is assessed using standard metrics for binary classification, including accuracy, precision, recall, and the $F1\rm{-}$score, defined as
@@
The normalized confusion matrix for the binary classification task, evaluated on the validation set and shown in Figure 1, demonstrates the strong performance of the trained RF classifier. The model correctly classifies 99.44% of quasars and 98.5% of stars, with minimal contamination between the two classes. This performance is further supported by an overall $F1\rm{-}$score of 0.9868, indicating consistently high precision and recall.
Figure 1: Normalized confusion matrix of the RF classifier computed on the validation set. The matrix is color-coded by the number of sources in each cell. Diagonal entries show the fraction of correctly classified objects (i.e., recall or completeness) for each class, while off-diagonal entries indicate the misclassification rates.
@@ IV Redshift Estimation for Quasar Candidates
We estimate photometric redshifts for the quasar candidates identified by the RF classifier using an RF-based regression model. Reliable redshift estimation is essential for characterizing the statistical properties of the candidate sample and for enabling subsequent population studies, particularly in regions where spectroscopic follow-up is incomplete.
@@
Photometric redshift estimation based solely on Gaia photometry is limited by the broad optical passbands and the lack of detailed spectral information. To improve regression performance, we incorporate additional optical photometry from the Pan-STARRS1 (PS1; PS1) survey, which provides five broadband filters ($g$, $r$, $i$, $z$, and $y$). For sources not covered by PS1, photometry from the NOIRLab Source Catalog DR2 (NSC DR2; NSC_DR2) is adopted as a substitute. These data enable the construction of more informative color indices that better trace redshift-dependent spectral features.
Figure 2: Distribution of RF-based photometric redshifts for the quasar candidates. The red histogram shows the full-sky candidate sample, and the blue histogram shows candidates at low Galactic latitude ($|b|<20^{\circ}$). The ordinate shows source counts (logarithmic scale).
In total, we adopt 18 input features for the RF regression model. These include the optical, infrared, and astrometric color indices grg-r, rir-i, izi-z, zyz-y, $z$-W1, $y$-W1, W1-W2, $G_{\rm BP}-G_{\rm RP}$, $_{\rm BP}-G, GG$$G_{\rm RP}$, $G_{\rm RP}$-W1, $G$-W1, and $G$-W2, as well as $\log(f_{x})$, $\log(f_{x}/f_{g})$, $\log(1+C^{*})$, $\log(1+z_{\mathrm{low}})$, and $\log(1+z_{\mathrm{up}})$. Here, $z_{\mathrm{low}}$ (redshift_qsoc_lower) and $z_{\mathrm{up}}$ (redshift_qsoc_upper) correspond to the 0.15866 and 0.84134 quantiles of the Gaia-based redshift estimate $z_{\mathrm{Gaia}}$, representing the lower and upper confidence bounds, respectively.
@@ V Results
V.1 Classification Results
We apply the trained RF classifier to the 22,434 previously unclassified sources in the CSC–WISE–GAIA sample, i.e., sources without reliable spectroscopic classifications. Of these, 7723 sources are classified as quasar candidates and 14,711 as stars. Figure 3 presents the distribution of the resulting quasar and stellar candidates in the $G$ versus $\mathrm{W1}-\mathrm{W2}$ color–magnitude diagram.
Figure 3: Distribution of the all-sky candidates in the $G$ versus W1-W2 diagram as classified by the RF model. Red and blue points denote quasar and stellar candidates, respectively; the top and right panels show the corresponding marginal distributions in $G$ and W1-W2. The $G$ magnitudes shown on the horizontal axis are corrected for Galactic extinction. We apply a uniform magnitude filter to all candidates, excluding sources with observed $G>21$ mag (prior to extinction correction) and removing very bright objects with $G<10$ mag.
Among the previously unclassified sources, 11,751 are located within the Galactic plane ($|b|<20^{\circ}$). Of these, 1110 sources (9.45%) are classified as quasar candidates, while the remaining 10,641 sources are classified as stars, reflecting the strong dominance of Galactic stellar populations at low Galactic latitudes.
To further suppress stellar contamination, we apply an additional selection based on Gaia proper motions. We adopt the probabilistic zero–proper-motion criterion introduced by Fu_2021; Fu_2024; Fu_2025, which explicitly accounts for measurement uncertainties. The probability density of zero proper motion, $f_{\rm PM0}$, is defined as
@@ V.1 Classification Results
where $x$ is the proper motion in R.A. (pmra), $y$ is the proper motion in Decl. (pmdec), $\sigma_{x}$ and $\sigma_{y}$ are the corresponding uncertainties, and $\rho$ is the correlation coefficient between $x$ and $y$ (pmra_pmdec_corr). For a given uncertainty level, sources with smaller proper motions yield higher values of $f_{\rm PM0}$ by construction.
Figure 4 shows the distributions of $\log(f_{\rm PM0})$ for spectroscopically confirmed quasars, spectroscopically confirmed stars, and the quasar candidates. We adopt a conservative threshold of $\log(f_{\rm PM0})\geq-4$, which efficiently removes stellar contaminants while retaining the majority of quasars. Applying this criterion excludes 1437 sources initially classified as quasar candidates.
Figure 4: Distributions of $\log(f_{\mathrm{PM}0})$, the probability density at zero proper motion derived from Gaia astrometry (see Eq. 8), for spectroscopically confirmed quasars (green), spectroscopically confirmed stars (yellow), and the GPQ candidates (white). The vertical dashed line marks the adopted threshold $\log(f_{\mathrm{PM}0})=-4$ used to suppress stellar contaminants; note that $f_{\mathrm{PM}0}$ is a probability density (not a probability) and therefore can exceed unity.
After the proper-motion filtering, the final quasar candidate sample contains 6286 sources across the entire sky. Among these, 863 sources are located behind the Galactic plane ($|b|<20^{\circ}$) and are identified as GPQ candidates. Of the GPQs, 514 sources have quasar membership probabilities greater than 0.8 and are therefore considered high-confidence GPQ candidates. The spatial distribution of the GPQ candidates in Galactic coordinates is shown in Figure 5. A subset of these high-confidence GPQ candidates is presented in Table 4. The whole selected GPQs and all-sky quasar candidates are available at https://github.com/jeepsmoker288/GPQs-and-all-sky-quasar-candidates.
@@ V.2 Properties of GPQ Candidates
Name R.A. Decl. W1 W2 $G$ log($f_{x}$) $P_{\rm QSO}$ $z_{\rm phot}$
@@
Note. All $G$ band magnitudes have been corrected for Galactic dust extinction using Planck dust map. log($f_{x}$) is the logarithm of X-ray flux in 0.5-7keV ($b$ band). $P_{\text{QSO}}$ is the classification probability of QSOs. $z_{\text{phot}}$ is the photometric redshift.
The GPQ candidates identified in this work exhibit distinct X-ray and multi-wavelength properties compared with the all-sky quasar candidate sample. In particular, GPQs show systematically harder X-ray spectra. Figure 6 compares the distributions of the hard–soft energy band hardness ratio (${\rm{HR}}_{hs}$) for quasars in the spectroscopic training set and for the quasar candidates, separated into all-sky and low Galactic latitude ($|b|<20^{\circ}$) subsamples. In both the training set and the candidate sample, sources located at $|b|<20^{\circ}$ display higher hardness ratios than all-sky sources.
Figure 5: Spatial distribution of GPQ candidates ($|b|<20^{\circ}$) in Galactic coordinates (Mollweide projection). Points are color-coded by the RF-derived quasar membership probability, $P_{\rm QSO}$ (color bar).
Figure 6: Normalized histograms of the $\rm{HR}_{hs}$ (hard – soft energy band hardness ratio, see table 3) for quasars. Filled histograms representing all-sky/galactic plane ($|b|<20^{\circ}$) training set samples, and step outlines for selected all-sky/galactic plane candidates.
@@
Figure 7 compares the distributions of optical ($G$), mid-infrared ($\rm{W1}$), and X-ray fluxes, as well as redshift (spectroscopic or RF-based photometric redshift), between the training-set quasars and the GPQ candidates. In all wavelength regimes, the GPQ candidates are systematically fainter than the quasars in the training set. The $G$ and $\rm{W1}$ magnitude distributions are shifted toward fainter values, and the X-ray flux distribution peaks at lower fluxes.
Figure 7: Normalized histograms of $G$, W1, log($f_{x}$) (logarithm of X-ray flux in 0.5-7keV) and redshift (here, the redshifts of the QSO candidates are estimated photometrically via RF.) for GPQs and GPQ candidates. Filled histograms representing training set samples, and step outlines for selected candidates.
These trends indicate that the present selection method is sensitive to quasars that lie beyond the effective depth of existing spectroscopic samples. The fainter optical and infrared magnitudes of the GPQ candidates are consistent with significant dust extinction in the Galactic plane, while their lower X-ray fluxes suggest that this work probes a previously underexplored population of faint or heavily obscured quasars behind the Galactic disk.
We further compare the GPQ candidates identified in this work with the unified all-sky quasar candidate catalog CatGlobe (Fu_2025). Of the 863 GPQ candidates, 617 have counterparts in CatGlobe, while 246 do not. The unmatched candidates show no significant differences from the matched subset in their color–magnitude distributions, X-ray fluxes, or redshift distributions, indicating that they likely belong to the same underlying quasar population. The presence of these additional candidates highlights the importance of incorporating X-ray information in quasar selection, particularly for uncovering quasars missed by purely optical and infrared approaches in the heavily obscured regions of the Galactic plane.
Figure 8: Spectra of the two identified GPQs from our pilot follow-up program with NGPS of Hale Telescope. We show flux- and wavelength-calibrated one-dimensional optical spectra of 2CXO J031404.4+403900 (catalog ) (upper) and 2CXO J031712.7+405301 (catalog ) (lower). The redshift inferred from identified emission lines is given in the upper-right corner of each panel. Strong emission features are labeled, and the spectra are displayed in the observed frame.

### v2

@@ III Method
III.1 The Random Forest Classifier
Figure 2: The flowchart of the machine learning pipeline used for GPQ candidate selection. The process begins with the construction of the multi-wavelength parent sample from Chandra, Gaia, and CatWISE. Crucially, to prevent data leakage, the labeled dataset ($N=13,724$) is split into a training set (80%) and a strictly independent test set (20%) before any hyperparameter tuning. The 5-fold cross-validation is performed exclusively within the training branch. The final Random Forest model is evaluated on the held-out test set before being applied to the unlabeled sources ($N=22,434$) to identify new candidates.
The overall workflow of our candidate selection pipeline is illustrated in Figure 2. We employ a RF classifier to distinguish quasars from stellar-type objects in the CSC–WISE–GAIA sample. RF is an ensemble learning algorithm based on decision trees and bootstrap aggregation, originally introduced by Ho (1995) and later formalized by Breiman (2001). By combining multiple weak learners trained on bootstrapped subsets of the data and randomly selecting input features at each split, the RF algorithm effectively mitigates overfitting and performs robustly in high-dimensional parameter space.
The RF classifier is implemented using the scikit-learn [^1] https://scikit-learn.org/stable/index.html machine-learning library (Pedregosa et al., 2011). Each decision tree is trained independently, and the final classification is determined by a majority vote across all trees. For each source, the fraction of trees voting for the quasar class is interpreted as the quasar membership probability, providing a quantitative measure of classification confidence.
To separate quasars from stellar-type objects, we adopt a total of 31 input features that incorporate multi-wavelength photometric, astrometric, and X-ray information. These features include X-ray fluxes and hardness ratios from CSC 2.1, optical and mid-infrared magnitudes and colors from Gaia DR3 and WISE, extinction-related quantities, and astrometric parameters such as proper motion and parallax significance. A complete list and description of the adopted features are provided in Table 3.
@@ III.1 The Random Forest Classifier
$\mathrm{HR}_{hm}$ ACIS hard (2.0-7.0 keV) - medium (1.2-2.0 keV) energy band hardness ratio
$\mathrm{HR}_{hs}$ ACIS hard (2.0-7.0 keV) - soft (0.5-1.2 keV) energy band hardness ratio
$\mathrm{HR}_{ms}$ ACIS medium (1.2-2.0 keV) - soft (0.5-1.2 keV) energy band hardness ratio
$N_{\mathrm{H,Gal}}$ Galactic $N_{H}$ column density in direction of source
$\log(f_{x})$ Logarithm of X-ray flux in 0.5-7.0 keV ($b$ band)
@@
$\log(f_{\mathrm{BP}})$ Logarithm of $BP$ band flux
$\log(f_{\mathrm{RP}})$ Logarithm of $RP$ band flux
@@
RPW1RP-W1 RPW1RP-W1 color
GW1G-W1 GW1G-W1 color
GW2G-W2 GW2G-W2 color
W1W2W1-W2 W1W2W1-W2 color
Missing values may arise from limited survey depth, nondetections, or incomplete coverage in specific bands. To allow uniform processing of incomplete data, all missing entries are assigned a sentinel value of 9999-9999. While large sentinel values can sometimes introduce artificial variance in distance-based models, Random Forest is naturally robust to them due to its tree-based architecture. The algorithm partitions the feature space using simple thresholds, effectively isolating sentinel values into independent branches without introducing artificial scaling issues. This approach preserves the physical information associated with missing data, allowing the model to handle such cases consistently during both training and prediction.
The RF model relies on several tunable hyperparameters whose optimization is essential for maximizing classification performance. To explicitly prevent data leakage, the complete labeled dataset ($N=13,724$) was first randomly partitioned into an 80% training/validation set and a 20% independent test set. To ensure an unbiased evaluation across the sky, this split was stratified by source type and regional category (i.e., within or outside the Galactic plane, defined as $|b|<20^{\circ}$).
We focused on three key parameters: the number of trees in the ensemble (n_estimators), the number of features considered at each split (max_features), and the maximum depth of individual trees (max_depth). Hyperparameter tuning was performed exclusively within the 80% training set using a grid search with stratified five-fold cross-validation. The optimal configuration was determined to be $\texttt{max\_features}=\log_{2}31$ and $\texttt{n\_estimators}=215$, while no explicit upper limit on max_depth was required based on the validation results.
III.2 Classification Performance and Feature Importance
Following parameter tuning, the final model was trained on the entire training set and evaluated on the untouched, strictly independent test set. Classification performance is assessed using standard metrics for binary classification, including accuracy, precision, recall, and $F1$-score, defined as:
@@ III.2 Classification Performance and Feature Importance
We first evaluate the performance of the trained RF classifier on the independent test set. As shown in the confusion matrices (Figure 3), the classifier achieves a high overall accuracy of 98.83%. However, considering the potential impact of the Galactic environment on the classification, we further assess the performance specifically at low Galactic latitudes ($|b|<20^{\circ}$). While the all-sky F1-score for quasars is 99.06%, it drops to 89.66% in the Galactic plane region. These results are summarized in Table 4. The slight decrease in performance at low Galactic latitudes is expected, given the effects of strong extinction and increased source confusion.
Figure 3: Normalized confusion matrices of the RF classifier evaluated on the independent test set for recall. The left panel shows the all-sky performance, while the right panel highlights the performance specifically within the Galactic plane ($|b|<20^{\circ}$). The labels “QSO” and “STAR” represent quasars and stellar-type objects, respectively. The color scale uses a logarithmic normalization to better visualize the classification of minority classes.
Table 4: Comparison of Classification Performance between All-Sky and Galactic Plane Samples
Region Class Precision (%) Recall (%) F1 (%)
All-Sky QSO 98.32 99.82 99.06
STAR 99.71 97.23 98.45
Galactic QSO 86.67 92.86 89.66
Plane STAR 99.87 99.73 99.80
Note.Metrics are calculated on the independent test set. The Galactic plane sample is defined as sources with $|b|<20^{\circ}$. “QSO” and “STAR” represent quasars and stellar-type objects.
To further understand the limitations of our binary classifier and address the potential contamination from specific stellar types, we analyzed the subclass composition of the 29 false positives (i.e., stellar objects misclassified as quasars) in the independent test set.
Although our training set is dominated by normal stars ($N=3452$) and YSOs ($N=1601$), it also includes small but representative samples of accreting sources, including high-mass X-ray binaries (HMXBs, $N=27$), low-mass X-ray binaries (LMXBs, $N=22$), and cataclysmic variables (CVs, $N=23$).
Among the 29 false positives, the vast majority ($N=27$, 93.1%) are normal stars. However, we also identified 2 X-ray binaries (1 HMXB and 1 LMXB) as contaminants. While XRBs constitute only $\sim 0.4\%$ of the training set, they account for 6.9% of the classification errors. This disproportionate representation confirms that accreting binaries, which share similar X-ray and optical properties with quasars, are indeed more challenging to distinguish.
In contrast, zero CVs were found among the false positives, indicating that our feature set effectively rejects these sources despite their accretion nature. Given the intrinsic rarity of XRBs compared to the vast population of field stars, their contribution to the total contamination remains critically low, justifying the effectiveness of our binary classification approach for the current sample size.
Figure 4: Relative importance of the input features for the random forest classifier. Features are ranked based on their mean decrease in Gini impurity across 10 independent runs with different random seeds, with error bars representing the standard deviation.
The feature importance analysis (Figure 4) reveals the physical drivers of our classification. The proper motion significance (PMSIG) emerges as the most critical feature, reflecting the fundamental astrometric difference between distant quasars, which are essentially stationary, and Galactic stars with detectable motions (e.g., Gaia Collaboration et al., 2023a). The mid-infrared color W1W2W1-W2 also ranks highly, as it effectively captures the characteristic infrared excess produced by hot dust in the vicinity of active galactic nuclei, a feature that distinguishes quasars from most stellar populations. Furthermore, the importance of the $RP$ magnitude suggests that in heavily obscured regions like the Galactic plane, longer-wavelength optical photometry is vital for detecting background sources through interstellar extinction.
@@ IV Redshift Estimation for Quasar Candidates
Applying the trained RF classifier to the 22,434 unclassified sources in the CSC–WISE–GAIA sample yields 7,681 quasar candidates. Within this sample, the 1,076 sources located at low Galactic latitudes ($|b|<20^{\circ}$) are defined as Galactic Plane Quasar (GPQ) candidates (see Section V.1 for formal selection and further filtering). In this section, we use an RF-based regression model to estimate photometric redshifts for these candidates. Reliable redshift estimation is essential for characterizing the statistical properties of the candidate sample and for enabling subsequent population studies, particularly in regions where spectroscopic follow-up is incomplete.
@@
Photometric redshift estimation based solely on Gaia photometry is limited by the broad optical passbands and the lack of detailed spectral information. To improve regression performance, we incorporate additional optical photometry from the Pan-STARRS1 (PS1; Chambers et al., 2016) survey, which provides five broadband filters ($g$, $r$, $i$, $z$, and $y$). For sources not covered by PS1, photometry from the NOIRLab Source Catalog DR2 (NSC DR2; Nidever et al., 2021) is adopted as a substitute. These data enable the construction of more informative color indices that better trace redshift-dependent spectral features.
In total, we adopt 18 input features for the RF regression model. These include the optical, infrared, and astrometric color indices grg-r, rir-i, izi-z, zyz-y, zW1z-W1, yW1y-W1, W1W2W1-W2, $G_{\mathrm{BP}}-G_{\mathrm{RP}}$, $_{\mathrm{BP}}-G, GG$$G_{\mathrm{RP}}$, $G_{\mathrm{RP}}-W1$, GW1G-W1, and GW2G-W2, as well as $\log(f_{x})$, $\log(f_{x}/f_{g})$, $\log(1+C^{*})$, $\log(1+z_{\mathrm{low}})$, and $\log(1+z_{\mathrm{up}})$. Here, $z_{\mathrm{low}}$ (redshift_qsoc_lower) and $z_{\mathrm{up}}$ (redshift_qsoc_upper) correspond to the 0.15866 and 0.84134 quantiles of the Gaia-based redshift estimate $z_{\mathrm{Gaia}}$, representing the lower and upper confidence bounds, respectively.
The final regression sample, constructed from the intersection of our labeled training set and the PS1/NSC photometric catalogs, consists of 7,332 spectroscopically confirmed quasars. The redshift distribution of this spectroscopic sample is shown as the green filled histogram in Figure 5, covering a wide range of $0<z<5$ which effectively spans the expected redshift space of our candidates.
We performed a quantitative evaluation of the regression model using the independent test set. The results are summarized in Figure 6. The model yields an RMSE of 0.3213 and a coefficient of determination ($R^{2}$) of 0.8384. These metrics, combined with the strong concentration of sources along the identity line in the density plot, demonstrate that our photometric redshift estimates are robust and free from significant latitude-dependent systematic biases.
Figure 5: Distribution of redshifts. The green filled histogram displays the spectroscopic redshift distribution of the training/test sample (N=N=7,332) used to build the regression model.
The red open histogram shows the RF-based photometric redshifts for the full-sky candidate sample, and the blue open histogram shows the candidates at low Galactic latitude ($|b|<20^{\circ}$). The ordinate shows source counts (logarithmic scale).
Figure 6: Quantitative evaluation of the regression model for photometric redshift estimation. The hexagonal binning plot displays the density of predicted redshifts ($z_{\mathrm{phot}}$) versus spectroscopic redshifts ($z_{\mathrm{spec}}$) for the independent test set. The logarithmic color scale indicates the source count per bin.
@@ V Results
V.1 GPQ candidate selection and additional filtering
We apply the trained RF classifier to the 22,434 previously unclassified sources in the CSC-WISE-GAIA sample, i.e., sources without reliable spectroscopic classifications. Of these, 7681 sources are classified as quasar candidates and 14,753 as stars. Figure 7 presents the distribution of the resulting quasar and stellar candidates in the $G$ versus W1W2W1-W2 color–magnitude diagram.
Among the previously unclassified sources, 11,751 are located within the Galactic plane ($|b|<20^{\circ}$). Of these, 1076 sources are classified as quasar candidates, while the remaining sources are classified as stellar-type objects, reflecting the strong dominance of Galactic stellar populations at low Galactic latitudes.
To further suppress stellar contamination, we apply an additional selection based on Gaia proper motions. We adopt the probabilistic zero–proper-motion criterion introduced by Fu et al. (2021, 2024, 2025), which explicitly accounts for measurement uncertainties. The probability density of zero proper motion, $f_{\mathrm{PM}0}$, is defined as
@@ V.1 GPQ candidate selection and additional filtering
where $x$ is the proper motion in R.A. (pmra), $y$ is the proper motion in Decl. (pmdec), $\sigma_{x}$ and $\sigma_{y}$ are the corresponding uncertainties, and $\rho$ is the correlation coefficient between $x$ and $y$ (pmra_pmdec_corr). For a given uncertainty level, sources with smaller proper motions yield higher values of $f_{\mathrm{PM}0}$ by construction.
Figure 7: Distribution of the all-sky candidates in the $G$ versus W1W2W1-W2 diagram as classified by the RF model. Red and blue points denote quasar and stellar-type candidates, respectively; the top and right panels show the corresponding marginal distributions in $G$ and W1W2W1-W2. The $G$ magnitudes shown on the horizontal axis are corrected for Galactic extinction. We apply a uniform magnitude filter to all candidates, excluding sources with observed $G>21$ mag (prior to extinction correction) and removing very bright objects with $G<10$ mag.
Figure 8 shows the distributions of $\log(f_{\mathrm{PM}0})$ for spectroscopically confirmed quasars, spectroscopically confirmed stellar-type objects, and the GPQ candidates. We adopt a conservative threshold of $\log(f_{\mathrm{PM}0})\geq-4$, which efficiently removes stellar contaminants while retaining the majority of quasars.
Figure 8: Distributions of $\log(f_{\mathrm{PM}0})$, the probability density at zero proper motion derived from Gaia astrometry (see Eq. 8), for spectroscopically confirmed quasars (green), spectroscopically confirmed stellar-type objects (yellow), and the GPQ candidates (white). The vertical dashed line marks the adopted threshold $\log(f_{\mathrm{PM}0})=-4$ used to suppress stellar contaminants; note that $f_{\mathrm{PM}0}$ is a probability density (not a probability) and therefore can exceed unity.
After the proper-motion filtering, the final quasar candidate sample contains 7570 sources across the entire sky. Among these, 1060 sources are located behind the Galactic plane ($|b|<20^{\circ}$) and are identified as GPQ candidates. Of the GPQs, 551 sources have quasar membership probabilities greater than 0.8 and are therefore considered high-confidence GPQ candidates. The spatial distribution of the GPQ candidates in Galactic coordinates is shown in Figure 9. A subset of these high-confidence GPQ candidates is presented in Table 5.
As mentioned in Section II.1, we evaluated the impact of source confusion in the Galactic plane using the probabilistic algorithm nway. Applying strict probability cuts with nway yielded 980 GPQ candidates, representing a relatively small reduction from the 1,060 candidates obtained via our nearest-neighbor approach. More importantly, we found that the strict nway procedure completely rejected two of our high-confidence candidates that were recently spectroscopically confirmed as quasars via the Next Generation Palomar Spectrograph (NGPS) on the Hale Telescope (see Section V.3). This observation directly confirms that while probabilistic matching filters out ambiguous blended sources, it penalizes completeness. Because our primary goal is to identify more quasar candidates in the Galactic plane, we retain the nearest-neighbor matching results.
@@ V.2 Properties of GPQ Candidates
Name R.A. Decl. $W1$ $W2$ $G$ $\log(f_{x})$ $P_{\mathrm{QSO}}$ $z_{\mathrm{phot}}$
@@
Note. All $G$ band magnitudes have been corrected for Galactic dust extinction using the Planck Collaboration et al. (2016) dust map. $\log(f_{x})$ is the logarithm of the X-ray flux in 0.5–7.0 keV ($b$ band). $P_{\mathrm{QSO}}$ is the classification probability of QSOs. $z_{\mathrm{phot}}$ is the photometric redshift.
The GPQ candidates identified in this work exhibit distinct X-ray and multi-wavelength properties compared with the all-sky quasar candidate sample. In particular, GPQs show systematically harder X-ray spectra. Figure 10 compares the distributions of the hard–soft energy band hardness ratio ($\mathrm{HR}_{hs}$) for quasars in the spectroscopic training set and for the quasar candidates, separated into all-sky and low Galactic latitude ($|b|<20^{\circ}$) subsamples. In both the training set and the candidate sample, sources located at $|b|<20^{\circ}$ display higher hardness ratios than all-sky sources.
Figure 9: Spatial distribution of GPQ candidates ($|b|<20^{\circ}$) in Galactic coordinates (Mollweide projection). Points are color-coded by the RF-derived quasar membership probability, $P_{\mathrm{QSO}}$ (color bar).
Figure 10: Normalized histograms of the $\mathrm{HR}_{hs}$ (hard – soft energy band hardness ratio, see Table 3) for quasars. Filled histograms represent all-sky/Galactic plane ($|b|<20^{\circ}$) training set samples, and step outlines represent selected all-sky/Galactic plane candidates.
@@
Figure 11 compares the distributions of optical ($G$), mid-infrared ($W1$), and X-ray fluxes, as well as redshift (spectroscopic or RF-based photometric redshift), between the training-set quasars and the GPQ candidates. In all wavelength regimes, the GPQ candidates are systematically fainter than the quasars in the training set. The $G$ and $W1$ magnitude distributions are shifted toward higher values, and the X-ray flux distribution peaks at lower fluxes.
Figure 11: Normalized histograms of $G$, $W1$, $\log(f_{x})$ (logarithm of X-ray flux in 0.5–7.0 keV), and redshift (here, the redshifts of the QSO candidates are estimated photometrically via RF) for GPQs and GPQ candidates. Filled histograms represent training set samples, and step outlines represent selected candidates.
These trends indicate that the present selection method is sensitive to quasars that lie beyond the effective depth of existing spectroscopic samples. The higher optical and infrared magnitudes of the GPQ candidates are consistent with significant dust extinction in the Galactic plane, while their lower X-ray fluxes suggest that this work probes a previously underexplored population of faint or heavily obscured quasars behind the Galactic disk.
We further compare the GPQ candidates identified in this work with the unified all-sky quasar candidate catalog CatGlobe (Fu et al., 2025). Of the 1060 GPQ candidates, 600 have counterparts in CatGlobe, while 460 do not. The unmatched candidates show no significant differences from the matched subset in their color–magnitude distributions, X-ray fluxes, or redshift distributions, indicating that they likely belong to the same underlying quasar population. The presence of these additional candidates highlights the importance of incorporating X-ray information in quasar selection, particularly for uncovering quasars missed by purely optical and infrared approaches in the heavily obscured regions of the Galactic plane.
Figure 12: Spectra of the two identified GPQs from our pilot follow-up program with the Next Generation Palomar Spectrograph (NGPS) on the Hale Telescope. We show flux- and wavelength-calibrated optical spectra of 2CXO J031404.4+403900 (catalog ) (upper panel) and 2CXO J031712.7+405301 (catalog ) (lower panel). The redshift inferred from identified emission lines is given in the upper-right corner of each panel. Strong emission features are labeled, and the spectra are displayed in the observed frame.